# Efficacy of Large Language Models in Systematic Reviews

### v1

@@ Abstract.
Large Language Models, Systematic Review, ESG, Sustainability, CSR, Financial Performance
conference: ; ; ccs: Computing methodologies Artificial intelligenceccs: General and reference Performanceccs: General and reference Reliabilityccs: Information systems Information retrievalccs: Applied computing Economicsccs: Applied computing Document management and text processing
1. Introduction
1.1. Background and Motivation
Environmental, Social, and Governance (ESG) investing has continued to grow, with nearly a 60% increase in the use of ESG as investment criteria within the last two decades (Umar et al., 2020). Increasing awareness of sustainability issues has made ESG considerations a mainstream component for asset managers and institutional investors. The United Nations Principles for Responsible Investment now has over 609 asset owners holding USD 121.3 trillion combined assets under management in 2021 (Kräussl et al., 2024). Additionally, the percentage of corporations releasing ESG reports has increased from 35% in 2010 to 86% in 2020 (Rouen et al., 2024).
However, there is no consensus on whether ”ESG pays.” Ľuboš Pástor, Robert F. Stambaugh, and Lucian A. Taylor [(2022)], for example, find that U.S. green stocks outperform brown stocks. In contrast, Juddoo et al. [(2023)] conclude impact investing strategies do not provide greater returns than traditional strategies. Many others have concluded that ESG does pay (Huynh and Xia, 2021; Hekman et al., 2022), that ESG doesn’t pay (Raghunandan and Rajgopal, 2022; Bolton and Kacperczyk, 2021), or that ESG has no impact on financial returns (Laplume et al., 2022; Amiraslani et al., 2023).
In addition to this dispute in the industry, there are concerns regarding the volume, transparency, and usefulness of ESG reports released by corporations (Burnaev et al., 2023). These inconsistencies complicate accurate measurement of a firm’s ESG performance and its impact on financial outcomes. Large Language Models (LLMs) offer a promising solution in this context. LLMs can rapidly interpret and analyze vast amounts of text and numerical information, enabling them to improve the efficiency and accuracy of ESG data analysis and serve as a tool to help companies familiarize themselves with ESG standards (Chopra et al., 2024). Thus, LLMs offer the potential to reduce the complexity of the ESG investing landscape (Kumar et al., 2022).
1.2. Problem Statement
@@ 2. Previous Works
In recent years, LLMs have made significant progress in their Natural Language Processing (NLP) abilities, marking an important milestone in real-world applications of Artificial Intelligence (AI). They can understand complex scenarios, analyze vast amounts of data, and generate relevant content, highlighting the immense potential present within their use in the sustainable financial industry (Chang et al., 2023; Wu et al., 2023). However, their development began with hand-coded models that relied on humans to determine sentiment.
Dev Shah et al. [(2018)] developed a manual sentiment analysis algorithm to predict the market movement of stocks within the pharmaceutical industry. They first gathered news articles on their stocks of interest and extracted the main text from those articles into n-grams (sets of words). Then, they created a data dictionary of 100 words covering a comprehensive list of topics specific to the pharmaceutical sector, classifying each word into a positive, negative, or neutral class of sentiment. Finally, the coded dictionary was used to classify the n-grams from the news text as ”positive,” ”neutral,” or ”negative,” corresponding to market predictions of stock price increase, remain steady, and decline, respectively. This method obtained a substantial prediction accuracy of 70.59%, highlighting the potential for sentiment analysis to play a significant role in increasing market efficiency (Shah et al., 2018). The researchers mention that a vital contributor to their high accuracy was the comprehensive data dictionary they made relying on domain expertise.
While Shah et al.’s [(2018)] model did not use machine learning techniques, Sandro Gössi et al. [(2023)] developed a more modern approach to the issue of sentiment analysis by fine-tuning the FinBERT model previously developed by Doug Araci [(2019)]. The FinBERT model is a fine-tuned version of the BERT language model (Devlin et al., 2019) for analyzing the sentiment of the financial text. Gössi et al. [(2023)] created a custom training set of 3535 complex sentences extracted from a dataset of Federal Open Market Committee (FOMC) financial texts, a significant increase in context from Shah et al. [(2018)]. Using this set to fine-tune the model, they saw significant increases in accuracy over the previous FinBert model and Shah et al.’s [(2018)] manual model, displaying the abilities of LLMs to extract meaning from nuanced language and financial jargon (Gössi et al., 2023). The performance benefits of continuous fine-tuning allude to the potential for future applications within niche areas such as ESG.
ESG scores are highly relevant to key players in the financial industry: asset owners determining investment risk, companies showcasing their environmental and social impacts, and regulators ensuring ethical standards are followed. ESG scoring is a meticulous process performed by key agencies, MSCI, Sustainalytics, ISS, Refinitiv, and FTSE Russell (Lykkesfeldt and Kjaergaard, 2022). To determine an overall ESG rating for a company, these agencies employ analysts to review historical data to understand drivers of ESG or hypothesize theoretical relationships that have not been tested. Each company determines its key metrics and corresponding weights to determine its own ESG score, leading to differences between the ratings of multiple agencies (Del Vitto et al., 2023). While this process has traditionally been performed by hand and statistical algorithms, LLMs can be a powerful tool to redefine ESG ratings (Zhao et al., 2024). Their capabilities of rapidly interpreting unstructured data, such as company reports, financial statements, and news articles, can significantly increase the efficiency of evaluating companies’ ESG performance. Furthermore, given unbiased training data, LLMs can remove human subjectivity and bias from all ratings, helping to reduce discrepancies across agencies (Zhao et al., 2024).
@@
Del Vitto et al. [(2023)] tested the efficacy of white-box algorithms such as Ridge and Lasso linear regression and black-box algorithms such as Random Forests and Artificial Neural Networks (ANN). The algorithms were tested on their ability to predict ESG scores given a data set comprising 13,052 firms that are stratified by industry sector and geographical region. The algorithms were designed to evaluate firms on 180 environmental, 218 social, and 152 governance attributes. Overall, the Ridge and Lasso algorithms performed the best with root mean squared errors (RMSE) as low as 0.071. Additionally, ANN algorithms outperformed the Ridge and Lasso algorithms within the Social pillar, with the lowest RMSE of 0.090 (Del Vitto et al., 2023).
3. Methods
@@
In this study, we assessed the performance of two base large language models (LLMs): Meta AI’s Llama 3 8B [^1] Accessed using the Groq API and OpenAI’s GPT-4 Omni (GPT-4o). Our evaluation was conducted on two sets of papers: 88 papers published between March 2020 and May 2024 and a previously compiled set of 238 papers from January 2015 to February 2020 by Atz et al.[(2021)]. Each LLM was tested using nine specific prompts in a 3x3 experimental design (Figure 5). Due to API token input limitations, Llama 3 8B was only provided with the abstracts of the papers, while GPT-4o was given the full PDF.
@@
3.1. Data Collection
We queried ProQuest, Science Direct, JSTOR, Google Scholar, and Social Science Research Network (SSRN) to find literature on ESG and financial performance published between March 2020 and May 2024. To ensure unbiased and high-quality papers were sourced, we filtered the results to include only papers from journals rated 3, 4, or 4* and classified as either ACCOUNT, FINANCE, ECON, ETHICS-CSR-MAN, or STRAT from the Chartered Association of Business 2021 Academic Journal Guide (Chartered Association of Business Schools, 2021). We compiled a corpus of 98 papers and hand-coded each in a typical systematic review format. However, we removed 10 papers, all from the Journal of Business Ethics, as they were deemed not salient enough (Atz et al., 2021). Of the remaining 88 papers, 14 manuscripts are not peer-reviewed. In addition to the new papers, we downloaded an older set from January 2015 to February 2020 comprising 238 papers from Atz et al. [(2021)].
3.2. Data Classification on New Set
The new data set labels followed the same structure as Atz et al.’s [(2023)] paper, as detailed in the online appendix (Atz et al., 2021). Three main questions were answered, and their responses were coded into the respective categories:
  1. (1)
@@ 3.2. Data Classification on New Set
3.3. Prompt Categories and Definitions
We split up three questions into three categories (1-3), which were structured incrementally with increasing context (A-C). The A-level prompts were basic and simply asked the questions and listed the possible classifications. The B-level prompts provided increased context from the Atz et al. [(2023)] codebook behind the meaning of the classifications, along with one or two examples. The C-level prompts included all of the information from the B-level and A-level but were designed to follow a chain-of-thought approach. They asked the model to take specific steps in determining its response. In addition, the model was asked to return a reasoning and confidence score for its prediction. This confidence score was derived within the prediction of the LLM and, therefore, should not be interpreted as a technical measure of confidence. These prompts were used for Llama 3 and base GPT-4o.
3.4. Custom GPT Development and Model Fine-tuning
The ”Custom GPT” model was developed using the ChatGPT interface’s ”GPTs” tool, where we provided a set of instructions and relevant attachments for the default model context. The three chain-of-through prompts (1C, 2C, and 3C) were combined into one large set of instructions for the ”Custom GPT.” This way, following the input of a paper, all three prompts would be answered within a single response and without additional prompting. Each paper was tested in a new chat session to ensure that previous interactions did not influence the model’s responses. In addition to a specialized prompt, the model was provided the old corpus 238 papers as a labeled training sample. [^2] Due to the design of OpenAI’s GPT builder, the old corpus served primarily as a reference rather than technical training data.
@@ 3.4. Custom GPT Development and Model Fine-tuning
3.5. Data Processing and Analysis
@@
(1) Weighted Power $\text{Weighted Power Mean}=\left(\frac{\sum w_{i}x_{i}^{2}}{\sum w_{i}}\right)^{\frac{1}{2}}$
4. Results
4.1. Old Data Set
Figure 1. Accuracy of interpretation of Llama 3 on prompts 1A-C and GPT-4o on prompts 1C
[…]
Results of LLMs on new data set prompt 1A through C
@@ 4.1. Old Data Set
4.2. New Data Set
Figure 2. Accuracy of interpretation of Llama 3 and GPT-4o on prompts 1A-C
[…]
Results of LLMs on new data set prompt 1A through C
Table 1. Accuracy of interpretation of Llama 3 and GPT-4o on prompts 1A-C overall and subsets
@@ 4.2. New Data Set
Table 2. Accuracy of interpretation of Llama 3 and GPT-4o on prompts 2A-C overall and subsets
@@
Table 3. Accuracy of interpretation of Llama 3 and GPT-4o on prompts 3A-C overall and subsets
@@
Table 4. Confidence score given by Llama 3 and GPT-4o on chain-of-thought prompts
@@
Figure 3. Base GPT-4o Misclassification Example on Prompt 1C
@@
Table 5. A-C Sub-prompt Agreement on Prompts 1-3
@@
Table 6. Weighted Power Mean of F1 score
@@
Table 6 shows the weighted power mean (with $p=2$) of the F1 scores for different LLM and prompt combinations across three scenarios.
Figure 4. Confusion Matrix for GPT4o (1A)
Figure 4 presents the confusion matrix for GPT-4o’s performance on the new data set with prompt 1A. ”Positive” samples were incorrectly classified as ”Mixed” 58% of the time. Overall, the model’s responses were skewed to the ”Mixed” category, incorrectly classifying true ”Positive” or ”Negative” studies as ”Mixed” frequently.
5. Discussion
5.1. Base Llama 3
@@
5.2. Base GPT-4o
Although base GPT-4o performed worse than Llama 3, it displayed highlights within certain prompts. For papers classified as ”Negative” in category 1, the model displayed relatively strong accuracies ranging from 82.6% to 87.0%. Because papers classified as ”Negative” are more likely to contain stronger language (need to find a source for this statement), the increase in accuracy may indicate that GPT-4o focuses more on searching for sentiment keywords throughout the paper. However, the model struggled with the ”Positive” class and misclassified 58% of ”Positive” papers as ”Mixed.” Furthermore, similar to Llama’s increase in performance for category 2, GPT-4o performed better on the ”Accounting-Based” paper, again highlighting the model’s ability to identify key information while ignoring the noise. Finally, GPT-4o was strong at classifying papers labeled as ”CSR” within category 3, which further supports the idea that GPT-4o can scan for keywords and sentiment well because often when CSR was used as a metric by the paper, it was explicitly stated as ”CSR” multiple times. The accuracy differences between GPT-4o and Llama 3 suggest a potential difference in training data between both LLMs. Additionally, the weighted power mean of the F1 score was the highest when using the chain-of-thought prompt, signifying that prompt design and structure play a significant role in the model outcome.
5.3. Fine-Tuned Models
@@
5.4. Evaluation of LLMs as a Whole
Ultimately, the application of LLMs within ESG shows potential. However, the base model LLMs are not completely reliable, with a maximum accuracy of 68.2% in the overall relationship between financial performance and ESG, 77.6% in the financial metrics category, and 81.6% in the sustainability metrics category. The base Llama 3 model consistently outperformed GPT-4o, yet it did not reach accuracy levels safe enough to replace human reviewers. To improve performance and accuracy, fine-tuning methods allow LLMs to serve as a tool to expedite traditional systematic reviews and transition to new best practices.
5.5. Future Works
Our research highlights the importance of selecting the appropriate prompt, model, and context to maximize the utility of LLMs. The chain-of-thought prompting style improved accuracy and provided insights into the model’s “thought process,” making it an area worth further exploration. Future research should focus on fine-tuning LLMs with larger and more diverse datasets to enhance their ability to interpret nuanced and complex information. With these better-equipped models, researchers can develop frameworks that integrate LLMs into existing systematic review workflows to assist in data extraction, synthesis, and reporting, ultimately streamlining the review process and maintaining the relevance of systematic reviews over time.
Acknowledgements.
We thank Ulrich Atz for the guidance and review and for providing the OpenAI API computing credits. We also thank the seminar participants at the NYU Stern Center for Sustainable Business for valuable feedback.
Appendix A Table and Figure Abbreviations

### v2

# Efficacy of Large Language Models for Systematic Reviews

@@ Abstract
Index Terms:
Large Language Models, Textual Analysis, ESG, Sustainability, CSR, Financial Performance
I Introduction
I-A Background and Motivation
Environmental, Social, and Governance (ESG) investing has continued to grow, with nearly a 60% increase in the use of ESG as investment criteria within the last two decades [1]. Increasing awareness of sustainability issues has made ESG considerations a mainstream component for asset managers and institutional investors. The United Nations Principles for Responsible Investment now has over 609 asset owners holding USD \$121.3 trillion combined assets under management in 2021 [2]. Additionally, the percentage of corporations releasing ESG reports has increased from 35% in 2010 to 86% in 2020 [3].
However, there is no consensus on whether ”ESG pays.” Ľuboš Pástor, Robert F. Stambaugh, and Lucian A. Taylor [4], for example, find that U.S. green stocks outperform brown stocks. In contrast, Juddoo et al. [5] conclude impact investing strategies do not provide greater returns than traditional strategies. Many others have concluded that ESG does pay [6, 7], that ESG doesn’t pay [8, 9], or that ESG has no impact on financial returns [10, 11].
In addition to this dispute in the industry, there are concerns regarding the volume, transparency, and usefulness of ESG reports released by corporations [12]. These inconsistencies complicate accurate measurement of a firm’s ESG performance and its impact on financial outcomes. Large Language Models (LLMs) offer a promising solution in this context. LLMs can rapidly interpret and analyze vast amounts of text and numerical information, enabling them to improve the efficiency and accuracy of ESG data analysis and serve as a tool to help companies familiarize themselves with ESG standards [13]. Thus, LLMs offer the potential to reduce the complexity of the ESG investing landscape [14].
I-B Problem Statement
@@ II Previous Works
In recent years, large language models (LLMs) have made significant progress in their Natural Language Processing (NLP) abilities, marking an important milestone in real-world applications of Artificial Intelligence (AI). These models can understand complex scenarios, analyze vast datasets, and generate relevant content, highlighting the immense potential within the sustainable financial industry [15, 16].
In the context of finance, Araci [17] developed FinBERT, a pre-trained language model specifically fine-tuned for financial sentiment analysis. Leveraging BERT’s transformer architecture, FinBERT was trained on a large corpus of financial texts to grasp the nuances of financial language, enabling it to effectively classify sentiments in financial news and reports. FinBERT outperformed state-of-the-art models at the time, with an accuracy increase of 15
Gössi et al. [18] then fine-tuned the FinBERT model on a custom training set of 3,535 complex sentences from Federal Open Market Committee (FOMC) financial texts. This fine-tuned model saw significant accuracy gains over the previous FinBERT model, demonstrating the ability of LLMs to extract meaning from nuanced language and financial jargon.
Furthermore, Sarmah et al. [19] investigated methods such as retrieval-augmented generation and the use of metadata to minimize hallucinations when extracting information from financial reports with LLMs. Meanwhile, Dolphin et al. [20] employed chain-of-thought prompting to enhance the LLM’s sentiment analysis performance and rigorously validated output to ensure dependable results, increasing reliable information retrieval from a broader array of news sources by over 400%.
In the specific context of this study—ESG analysis—AI models have been instrumental in redefining methods of scoring ESG performance for companies. Their effectiveness in interpreting financial data makes them powerful tools for analyzing vast amounts of unstructured data, such as company reports, financial statements, and news articles, to predict a company’s ESG performance [21]. Moreover, if AI models are trained on unbiased data, they can help eliminate human subjectivity and bias from ratings, reducing discrepancies across agencies [21].
@@
Del Vitto et al. [22] evaluated white-box algorithms (Ridge and Lasso linear regression) and black-box algorithms (Random Forests and Artificial Neural Networks) on their ability to predict ESG scores using a training set of 13,052 firms stratified by industry sector and geographic region. Ridge and Lasso algorithms achieved root mean squared errors (RMSE) as low as 0.071, while artificial neural networks (ANNs) outperformed Ridge and Lasso within the Social pillar, achieving a lowest RMSE of 0.090 [22].
III Methods
@@
In this study, we assessed the performance of two base large language models (LLMs): Meta AI’s Llama 3 8B [^1] Accessed using the Groq API and OpenAI’s GPT-4 Omni (GPT-4o). Our evaluation was conducted on two sets of papers: 88 papers published between March 2020 and May 2024 and a previously compiled set of 238 papers from January 2015 to February 2020 by Atz et al. [23]. Each LLM was tested using nine specific prompts in a 3x3 experimental design (Fig. 5). Due to API token input limitations, Llama 3 8B was only provided with the abstracts of the papers, while GPT-4o was given the full PDF.
@@
III-A Data Collection
We queried ProQuest, Science Direct, JSTOR, Google Scholar, and Social Science Research Network (SSRN) to find literature on ESG and financial performance published between March 2020 and May 2024. To ensure unbiased and high-quality papers were sourced, we filtered the results to include only papers from journals rated 3, 4, or 4* and classified as either ACCOUNT, FINANCE, ECON, ETHICS-CSR-MAN, or STRAT from the Chartered Association of Business 2021 Academic Journal Guide [24]. We compiled a corpus of 98 papers and hand-coded each in a typical systematic review format. However, we removed 10 papers, all from the Journal of Business Ethics, as they were deemed not salient enough [23]. Of the remaining 88 papers, 14 manuscripts are not peer-reviewed. In addition to the new papers, we downloaded an older set from January 2015 to February 2020 comprising 238 papers from Atz et al. [23].
III-B Data Classification on New Set
The new data set labels followed the same structure as Atz et al.’s [25] paper, as detailed in the online appendix [23]. Three main questions were answered, and their responses were coded into the respective categories:
  1. 1.
@@ III-B Data Classification on New Set
III-C Prompt Categories and Definitions
We split up three questions into three categories (1-3), which were structured incrementally with increasing context (A-C). The A-level prompts were basic and simply asked the questions and listed the possible classifications. The B-level prompts provided increased context from the Atz et al. [25] codebook behind the meaning of the classifications, along with one or two examples. The C-level prompts included all of the information from the B-level and A-level but were designed to follow a chain-of-thought approach. They asked the model to take specific steps in determining its response. In addition, the model was asked to return a reasoning and confidence score for its prediction. This confidence score was derived within the prediction of the LLM and, therefore, should not be interpreted as a technical measure of confidence. These prompts were used for Llama 3 and base GPT-4o.
III-D Custom GPT Development and Model Fine-tuning
The ”Custom GPT” model was developed using the ChatGPT interface’s ”GPTs” tool, where we provided a set of instructions and relevant attachments for the default model context. The three chain-of-through prompts (1C, 2C, and 3C) were combined into one large set of instructions for the ”Custom GPT.” This way, following the input of a paper, all three prompts would be answered within a single response and without additional prompting. Each paper was tested in a new chat session to ensure that previous interactions did not influence the model’s responses. In addition to a specialized prompt, the model was provided the old corpus 238 papers as a labeled training sample. [^2] The old corpus served primarily as a reference rather than modifying the model’s parameters.
@@ III-D Custom GPT Development and Model Fine-tuning
III-E Data Processing and Analysis
@@
Weighted Power $\text{Weighted Power Mean}=\left(\frac{\sum w_{i}x_{i}^{2}}{\sum w_{i}}\right)^{\frac{1}{2}}$ (1)
IV Results
IV-A Old Data Set
Figure 1: Accuracy of interpretation of Llama 3 on prompts 1A-C and GPT-4o on prompt 1A
@@ IV-A Old Data Set
IV-B New Data Set
Figure 2: Accuracy of interpretation of Llama 3 and GPT-4o on prompts 1A-C.
TABLE I: Accuracy of interpretation of Llama 3 and GPT-4o on prompts 1A-C overall and subsets
@@ IV-B New Data Set
TABLE II: Accuracy of interpretation of Llama 3 and GPT-4o on prompts 2A-C overall and subsets
@@
TABLE III: Accuracy of interpretation of Llama 3 and GPT-4o on prompts 3A-C overall and subsets
@@
TABLE IV: Confidence score given by Llama 3 and GPT-4o on chain-of-thought prompts
@@
Figure 3: Base GPT-4o Misclassification Example on Prompt 1C
@@
TABLE V: A-C Sub-prompt Agreement on Prompts 1-3
@@
TABLE VI: Weighted Power Mean of F1 score
@@
Table VI shows the weighted power mean (with $p=2$) of the F1 scores for different LLM and prompt combinations across three scenarios.
Figure 4: Confusion Matrix for GPT4o (1A)
Fig. 4 presents the confusion matrix for GPT-4o’s performance on the new data set with prompt 1A. ”Positive” samples were incorrectly classified as ”Mixed” 58% of the time. Overall, the model’s responses were skewed to the ”Mixed” category, incorrectly classifying true ”Positive” or ”Negative” studies as ”Mixed” frequently.
V Discussion
V-A Base Llama 3
@@
V-B Base GPT-4o
Although base GPT-4o performed worse than Llama 3, it displayed highlights within certain prompts. For papers classified as ”Negative” in category 1, the model displayed relatively strong accuracies ranging from 82.6% to 87.0%. Because papers classified as ”Negative” often state this explicitly, the increase in accuracy may indicate that GPT-4o focuses more on searching for sentiment keywords throughout the paper. However, the model struggled with the ”Positive” class and misclassified 58% of ”Positive” papers as ”Mixed.” Furthermore, similar to Llama’s increase in performance for category 2, GPT-4o performed better on the ”Accounting-Based” paper, again highlighting the model’s ability to identify key information while ignoring the noise. Finally, GPT-4o was strong at classifying papers labeled as ”CSR” within category 3, which further supports the idea that GPT-4o can scan for keywords and sentiment well because often when CSR was used as a metric by the paper, it was explicitly stated as ”CSR” multiple times. The accuracy differences between GPT-4o and Llama 3 suggest a potential difference in training data between both LLMs. Additionally, the weighted power mean of the F1 score was the highest when using the chain-of-thought prompt, signifying that prompt design and structure play a significant role in the model outcome.
V-C Fine-Tuned Models
@@
V-D Evaluation of LLMs as a Whole
Ultimately, the application of LLMs within ESG shows potential. However, while LLMs can analyze vast amounts of data, the base model accuracy levels indicate that there is still room for improvement in terms of accuracy and deeper understanding. Specifically, the maximum accuracy achieved by base LLM models in determining the relationship between financial performance and ESG metrics was 68.2%. When broken down into specific categories, models reached an accuracy of 77.6% for financial metrics and 81.6% for sustainability metrics. Notably, the base Llama 3 model consistently outperformed GPT-4o across these tasks, yet even Llama 3’s performance did not achieve an accuracy threshold that could replace human reviewers. This gap suggests that LLMs in their current form can support, but not fully automate, ESG analysis workflows.
To address these challenges, fine-tuning methods have been identified as a viable approach to enhance LLM accuracy and reliability. By training LLMs on ESG-specific datasets or tailoring them to a particular set of review standards, models can be refined to more accurately interpret and evaluate complex ESG data. With such fine-tuning, LLMs can potentially act as tools to expedite traditional systematic reviews, offering initial insights and analyses that streamline the review process while enabling human experts to focus on higher-level evaluation and decision-making. This approach positions LLMs as complementary assets to human reviewers rather than replacements, bridging the gap between automation and expert analysis in ESG-related applications.
V-E Future Works
Our research underlines the importance of choosing the appropriate model architecture, prompting techniques, and contextual information to maximize the effectiveness of LLMs in ESG analysis. Notably, chain-of-thought prompting emerged as an effective method, enhancing accuracy by guiding the model through logical steps and clarifying its “thought process.” This approach provided greater transparency into the model’s reasoning, making chain-of-thought prompting a valuable area for further exploration and refinement in ESG applications.
Future research should prioritize the fine-tuning of LLMs using larger, more diverse, and ESG-focused datasets to improve their capability to interpret nuanced and complex information. Additionally, efforts should be directed toward developing frameworks that seamlessly integrate LLMs into existing systematic review workflows. With better-equipped models, researchers can employ LLMs to assist in essential review tasks, such as data extraction, synthesis, and reporting. These enhanced models have the potential to streamline the systematic review process while improving consistency and efficiency, ultimately contributing to the ongoing relevance and accuracy of ESG-related reviews and ensuring that findings remain aligned with industry standards over time.
VI Abbreviations and Acronyms